A credit union has issued a 2 years loan of $6,000. Simple interest is charged 

at a rate of 8 percent per year. The principal plus interest is to be repaid at the 

end of second year. Compute the interest for the 2 – year period. What amount 

will be repaid at the end of the second year.   


1
Expert's answer
2021-04-20T15:58:51-0400

Interest ="\frac{PRT}{100}"

Where T is the time =2 years, P is the principal = $6000, and R is the rate at 8% (0.08)


Interest= "6000\u00d70.08\u00d72"

= "960"

The interest to be paid at the end of the 2-year period is $960


The amount to be repaid at the end of 2 years is given as

"FV=P(1+rn)"

Where FV is the future amount to be paid, R is the rate, and n is the number of years to be used in repaying the loan

"FV=6000(1+(0.08\u00d72))"

"=6000(1.16)"

"=6960"

The final amount to be repaid after 2 years is $6960